OpenStudy (anonymous):

Which equation is related to this equation? w divide 9 = 14

OpenStudy (anonymous):

Is this multiple choice or are you just being asked to write an equivalent equation?

OpenStudy (anonymous):

multiple choice

OpenStudy (anonymous):

A. w= 14 divide 9 B. w= 14 x 9 C. w= 14 - 9 D. w= 14 + 9

OpenStudy (anonymous):

Alright, well, know that using the operators multiplication and division, as long as you perform the same operation on all parts of the equation, you will arrive at an equivalent equation.